# Long-Context Generalization with Sparse Attention

### v2

@@ Abstract
Transformer-based architectures traditionally employ softmax to compute attention weights, which produces dense distributions over all tokens in a sequence. While effective in many settings, this density has been shown to be detrimental for tasks that demand precise focus on fixed-size patterns: as sequence length increases, non-informative tokens accumulate attention probability mass, leading to dispersion and representational collapse. We show in this paper that sparse attention mechanisms using $\alpha$-entmax can avoid these issues, due to their ability to assign exact zeros to irrelevant tokens. Furthermore, we introduce Adaptive-Scalable Entmax (ASEntmax), which endows $\alpha$-entmax with a learnable temperature parameter, allowing the attention distribution to interpolate between sparse (pattern-focused) and dense (softmax-like) regimes. Finally, we show that the ability to locate and generalize fixed-size patterns can be further improved through a careful design of position encodings, which impacts both dense and sparse attention methods. By integrating ASEntmax into standard transformer layers alongside proper positional encodings, we show that our models greatly outperform softmax, scalable softmax, and fixed-temperature $\alpha$-entmax baselines on long-context generalization.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has become the foundation of modern large language models (LLMs), establishing new benchmarks across diverse domains. However, as researchers push these models toward increasingly longer contexts—from thousands to millions of tokens—several fundamental limitations emerge that can be traced to the softmax transformation used in attention. Three critical limitations stand out: representational collapse occurs due to softmax’s inability to maintain distinct attention patterns as sequence length grows, erasing meaningful distinctions between tokens (Barbero et al., 2024); over-squashing is exacerbated by softmax’s dense probability distribution, leading to exponential dilution of gradients (Alon & Yahav, 2021; Barbero et al., 2024); and attention dispersion arises from softmax’s fundamental property that forces probability mass to be distributed across all tokens, with attention weights necessarily approaching an uniform distribution as context grows (Veličković et al., 2024; Nakanishi, 2025).
@@
In this paper, we address the root cause of these problems by replacing softmax with $\alpha$-entmax (Peters et al., 2019), a differentiable sparse transformation that induces probability distributions where irrelevant tokens receive exactly zero attention. While $\alpha$-entmax has been used successfully in transformers (Correia et al., 2019), its length generalization properties, to the best of our knowledge, have never been studied. We show theoretically and empirically that $\alpha$-entmax consistently helps to address challenges in long context modeling. Our key contributions include: [^1] Our code is available at https://github.com/deep-spin/adasplash.
@@
  • Representational preservation: We demonstrate that $\alpha$-entmax is able to prevent representational collapse by maintaining distinguishable token representations even in extremely long contexts.
  • Over-squashing alleviation: We show how $\alpha$-entmax reduces the number of gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, where $L$ is the number of layers, significantly strengthening gradient flow for long-range dependencies.
  • Adaptive-scalable $\alpha$-entmax: We introduce ASEntmax, which adaptively adjust sparsity based on sequence length to maintain optimal token selection even in extremely long contexts. Furthermore, we develop NAPE (NoPE + ALiBi Positional Encoding), a hybrid scheme that combines attention heads between content-focused processing and locality-aware positional encoding, leading to a synergistic interaction with $\alpha$-entmax.
  • Empirical validation: We demonstrate substantial performance improvements across diverse tasks. For example, as shown in Figure 1, ASEntmax achieves 76.7% accuracy on associative recall at 65K tokens after training on just 64 tokens—a 1000$\times$ length extrapolation.
Figure 1: Comparison of softmax and $\alpha$-entmax for long-context length on Multi-query Multi-token Associative Recall (left) and Reverse task (right). SSMax represents the Scalable Softmax approach by Nakanishi (2025). While all methods benefit from using NAPE, our adaptive-scaling version of $\alpha$-entmax exhibits the best extrapolation results, effectively handling extremely long sequences.
@@ 2.2 $\alpha$-entmax
where $[\cdot]_{+}:=\max(0,\cdot)$ and $\tau:\mathbb{R}^{n}\to\mathbb{R}$ yields a threshold that ensures the resulting distribution sums to $1$. A key property of $\alpha$-entmax is that tokens with scores below the threshold receive exactly zero probability, creating sparse attention patterns. When $\alpha\rightarrow 1^{+}$, this reduces to the standard softmax function. The sparsity level increases with $\alpha$, with $\alpha=2$ corresponding to the sparsemax function (Martins & Astudillo, 2016). Figure 2 illustrates $\alpha$-entmax for different values of $\alpha$. We provide more information on $\alpha$-entmax in §A. While $\alpha$-entmax is a suitable choice for sparse attention, its theoretical and empirical impact on long inputs is still unclear. In the next section, we demonstrate how it fundamentally changes the way attention behaves for long contexts.
@@ 3.1 Non-Vanishing Attention Probabilities
Furthermore, for $\alpha>1$, the difference $\alpha\text{-entmax}(\bm{x})_{n}-\alpha\text{-entmax}(\bm{x}^{*})_{n+1}$ can take any value in $[0,\alpha\text{-entmax}(\bm{x})_{n}]$ by appropriate choice of $b$.
The proof can be found in §C.1. This result demonstrates a fundamental difference between softmax and $\alpha$-entmax. Unlike softmax, where adding a new token always reduces the attention probability of existing tokens strictly, $\alpha$-entmax allows a distinct behavior: the attention probability can remain unchanged. This occurs because the $\alpha$-entmax’s thresholding effect allows tokens with logits below a certain threshold to receive exactly zero attention, letting the model focus only on the relevant tokens.
@@ 3.2 Attention Dispersion and Concentration
    This is impossible for $\alpha=1$ (softmax), for which we always have $\operatorname{softmax}(\bm{z})_{i}<\operatorname{softmax}(\bm{z}^{*})_{i}$.
@@
    $\lim_{n\to\infty}\frac{H(\operatorname{softmax}(\bm{z}_{1:n}/\theta))}{\log n}=1.$ (6)
@@
The full proof can be consulted in §D. The key takeaway of this result is that the entropy of attention distributions reveals how concentrated or dispersed they are across tokens. While softmax distributions approach maximum entropy $\Theta(\log n)$ as the sequence length increases (indicating complete dispersion), $\alpha$-entmax distributions maintain bounded entropy $\mathcal{O}(\log s)$ where $s$ is the support size. This allows models with $\alpha$-entmax to maintain focused, low-entropy attention patterns even when processing extremely long sequences, as long as the support size is smaller than the full sequence length. This non-dispersion property means transformers with $\alpha$-entmax can scale to very long contexts without the attention becoming dispersed, maintaining their ability to focus on relevant information regardless of how much additional context is present. However, attention dispersion is not the only obstacle to effective long-sequence modeling.
@@ 3.3 Representational Preservation and Over-squashing Alleviation
Two other critical challenges in long-context transformers are representational collapse and over-squashing—both fundamentally linked to the properties of softmax attention (Barbero et al., 2024). To clarify these concepts, we provide precise definitions below.
{definition}
[Representational Collapse and Over-squashing] A transformer model suffers from representational collapse when token representations become similar as sequence length grows: for any $\epsilon>0$, there exists $n_{\epsilon}$ such that $\|\bm{v}_{i}^{(L)}-\bm{v}_{j}^{(L)}\|_{1}<\epsilon$ for all $i,j\leq n$ when $n>n_{\epsilon}$.
Over-squashing occurs when gradient signals are exponentially diluted as they flow through $L$ layers, with $\|\frac{\partial\bm{y}_{n}}{\partial\bm{v}_{i}^{(0)}}\|<\epsilon$ for any $\epsilon>0$ due to the $\mathcal{O}(n^{L})$ gradient paths in softmax attention.
Building on these definitions, we show in §C.2 and §C.3 that $\alpha$-entmax addresses both issues: {proposition}[Representational Preservation] There are choices of $\bm{v}^{(0)}\in\mathbb{R}^{n\times d}$ and $\bm{v}^{*(0)}\in\mathbb{R}^{(n+1)\times d}$ such that $\alpha$-entmax ($\alpha>1$) can maintain distinct token representations even as $n\to\infty$: $\|\bm{v}_{n}^{(L)}-\bm{v}_{n+1}^{*(L)}\|_{1}\geq c$ for any constant $c>0$.
@@
[Over-squashing Alleviation] With support size $s\ll n$, $\alpha$-entmax with $\alpha>1$ reduces gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, alleviating over-squashing via stronger gradient signals.
To empirically validate these properties, we conducted controlled experiments with transformer architectures employing different attention mechanisms, fully described in §C. For representational collapse, we measured how token representations change when sequence length increases. For over-squashing, we analyzed gradient flow through an 8-layer network on a copying task, where the model must copy information across long distances. The results are shown in Figure 3. The left panel shows the $L_{1}$ norm of representation differences between original and extended sequences after multiple transformer layers. With softmax, this difference rapidly approaches zero as sequence length increases. In contrast, $\alpha$-entmax maintains bounded differences even at extreme lengths (128K tokens), with higher $\alpha$ values providing stronger preservation. The right panel illustrates gradient flow strength across sequence lengths, showing how $\alpha$-entmax maintains significantly stronger gradient signals across layers compared to softmax.
Figure 3: Empirical verification of representational collapse and over-squashing. Left: Representation difference ($L_{1}$ norm) after 6 layers across sequence lengths. Softmax ($\alpha=1.0$) exhibits inevitable collapse, while $\alpha$-entmax maintains bounded differences even at extreme lengths. Right: Gradient norms across layers in a transformer ($n=256$). $\alpha$-entmax maintains up to 6×6$\times$6 × stronger gradient signals than softmax, particularly in earlier layers, demonstrating its ability to alleviate over-squashing.
@@ 4.1 Controlling Sparsity in Long Contexts via ASEntmax
We start by introducing the following remark for a sequence of Gaussian random variables.
{remark}
[Normal Concentration] Let the attention logits $\bm{z}=(z_{1},\dots,z_{n})$ be IID $\mathcal{N}(0,\sigma)$ random variables and denote
$M:=\max_{1\leq i\leq n}z_{i},\qquad m:=\min_{1\leq i\leq n}z_{i},\qquad\Delta:=M-m.$ (8)
By standard results on the asymptotic behavior of extreme values of normal random variables (Kamath, 2015), the expected range satisfies $\mathbb{E}[\Delta]\leq 2\sigma\sqrt{2\log n}$ and $\lim_{n\rightarrow\infty}\frac{\mathbb{E}[\Delta]}{2\sigma\sqrt{2\log n}}=1.$
Since the logit spread increases only as $\sqrt{\log n}$, a fixed temperature can cause excessive sparsity in long contexts. To address this, we propose Adaptive-Scalable $\alpha$-entmax (ASEntmax), which adaptively scales the temperature for each attention head and query based on the context size and content:
@@
where $\bm{w}_{\beta},\bm{w}_{\gamma}\in\mathbb{R}^{d}$ are learnable, head-specific projection vectors. The scaling factor $(\log n)^{\gamma}$ in Eq. 9 is directly motivated by the asymptotic behavior of normal logits from Remark 4.1. This allows the model learn a slowly rising ($\gamma>0$) or dampening ($\gamma<0$) temperature schedule without interfering in the positional encodings (which would happen with negative values of $\beta$). Specifically, for IID Gaussian logits $\mathcal{N}(0,\sigma)$, when $\delta=0$ and $\gamma=-0.5$ the scaling counteracts the growth of logit ranges ($\Delta_{n}$) as context increases:
@@
Figure 4: Learned positions per head. Besides a simple linear fit baseline ($\beta n$), we also show the fit of a log model ($\beta\log n$) and a log-power model ($\beta(\log n)^{\gamma}$), inherited by ASEntmax.
@@
To empirically validate the importance of the parameter $\gamma$ in our proposed scaling formulation, we conducted experiments on a language modeling task using a 120M-parameter transformer trained on 5B tokens from the FineWeb dataset (Penedo et al., 2024). Following the methodology of Nakanishi (2025), we implemented learnable scaling parameters for the attention logits, but with a key difference: while they used a global scaling parameter, we learn separate scaling parameters for each attention head, motivated by Correia et al. (2019)’s finding that attention heads develop distinct sparsity patterns.
Figure 4 presents the learned scaling behaviors for representative attention heads, along with fitted curves from different scaling models. The results clearly demonstrate that scaling parameters vary significantly across heads, and that a simple log-scaling model $\beta\log n$ (as used by Nakanishi (2025)) provides poor fits for many heads. In contrast, our proposed form $\delta+\beta(\log n)^{\gamma}$ provides consistently better fits capturing the diverse scaling behaviors across different attention heads. The complete distribution of fitted $\beta$ and $\gamma$ values across all heads is provided in §F, and additional training details can be found in §F.1.
4.2 Interacting with Positional Encoding
Positional bias is known to affect attention dynamics and generalization (Barbero et al., 2025; Wu et al., 2025). In this work, we also introduce NAPE (NoPE + ALiBi Positional Encoding), a hybrid scheme in which half of the attention heads receive no positional encoding (NoPE), enabling flexible, data-driven position inference, while the other half use a fixed monotonic bias (ALiBi) to encourage locality. This complementary design allows some heads to focus on global patterns while others encode recency, combining the strengths of both.
While NAPE is not central to our theoretical results, we find it plays a crucial empirical role in enabling robust length generalization, especially when paired with the sparse inductive bias of $\alpha$-attention. As detailed in our theoretical and empirical analysis in §E, the NoPE component of NAPE enables models to develop effective relative positional encoding—supporting the flexibility hypothesis of Kazemnejad et al. (2023)—and, more broadly, $\alpha$-entmax interacts with positional encodings in ways that fundamentally alter attention behavior compared to softmax, such as creating hard attention windows with ALiBi or inducing frequency-dependent patterns with RoPE.
@@ 5 Experiments
All synthetic tasks are trained with a decoder-like transformer. Moreover, we evaluate models in extreme settings by using as few layers as possible as our goal is to test the attention mechanism coupled with our positional-encoding strategy, rather than the scaling capabilities of transformer. However, for the Reverse task, we increment the layer count until the softmax baseline generalizes to at least 1.5$\times$ the in-distribution length, as this task proved especially challenging for softmax models. To improve length extrapolation in models employing RoPE, we apply a RoPE scaling factor of 16. For NAPE, we set ALiBi slopes to $m=\frac{1}{h}$, where $h$ is the head index. Finally, we use $\delta=1$ in ASEntmax models. More experimental details are described in §G.

### v3

@@ Abstract
[…]
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has become the foundation of modern large language models (LLMs), establishing new benchmarks across diverse domains. However, as researchers push these models toward increasingly longer contexts—from thousands to millions of tokens—several fundamental limitations emerge that can be traced to the softmax transformation used in attention. Three critical limitations stand out: representational collapse occurs due to softmax’s inability to maintain distinct attention patterns as sequence length grows, erasing meaningful distinctions between tokens (Barbero et al., 2024); over-squashing is exacerbated by softmax’s dense probability distribution, leading to exponential dilution of gradients (Alon & Yahav, 2021; Barbero et al., 2024); and attention dispersion arises from softmax’s fundamental property that forces probability mass to be distributed across all tokens, with attention weights necessarily approaching an uniform distribution as context grows (Veličković et al., 2025; Nakanishi, 2025).
@@
In this paper, we address the root cause of these problems by replacing softmax with $\alpha$-entmax (Peters et al., 2019), a differentiable sparse transformation that induces probability distributions where irrelevant tokens receive exactly zero attention. While $\alpha$-entmax has been used successfully in transformers (Correia et al., 2019; Gonçalves et al., 2025), its length generalization properties, to the best of our knowledge, have never been studied. We show theoretically and empirically that $\alpha$-entmax consistently helps to address challenges in long context modeling.
@@
  • Representational preservation: $\alpha$-entmax attention with sparse support can avoid representational collapse and reduces the number of gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, alleviating over-squashing by strengthening the gradient flow for long-range dependencies.
  • Adaptive-scalable $\alpha$-entmax: We introduce ASEntmax, which adaptively adjust sparsity based on sequence length to maintain optimal token selection even in extremely long contexts.
[…]
Our code is available at https://github.com/deep-spin/asentmax.
Figure 1: Long-context generalization on Multi-query Multi-token Associative Recall (left) and Max Retrieval (right). SSMax represents the Scalable Softmax approach by Nakanishi (2025), and Adaptive Temperature (Adapt. Temp.) represents the approach by Veličković et al. (2025). While all methods benefit from using NAPE (NoPE + ALiBi), our adaptive-scaling version of $\alpha$-entmax exhibits the best extrapolation results, effectively handling extremely long sequences.
@@ 2.2 $\alpha$-entmax
where $\cdot]_{+}:=\max(0,\cdot)$ and $\tau:\mathbb{R}^{n}\to\mathbb{R}$ yields a threshold that ensures the resulting distribution sums to $1$. A key property of $\alpha$-entmax is that tokens with scores below the threshold receive exactly zero probability, creating sparse attention patterns. When $\alpha\rightarrow 1^{+}$, this reduces to the standard softmax function. The sparsity level increases with $\alpha$, with $\alpha=2$ corresponding to the sparsemax function (Martins & Astudillo, 2016). Figure 2 illustrates $\alpha$-entmax for different values of $\alpha$. We provide more information on $\alpha$-entmax in §A. While $\alpha$-entmax is a suitable choice for sparse attention, its theoretical and empirical impact on long inputs is still unclear. In the next section, we demonstrate how it fundamentally changes the way attention behaves for long contexts.
@@ 3.1 Non-Vanishing Attention Probabilities
Furthermore, for $\alpha>1$, the difference $\alpha\text{-entmax}(\bm{x})_{n}-\alpha\text{-entmax}(\bm{x}^{*})_{n+1}$ can take any value in $0,\alpha\text{-entmax}(\bm{x})_{n}]$ by appropriate choice of $b$.
The proof can be found in §D.1. This result demonstrates a fundamental difference between softmax and $\alpha$-entmax. Unlike softmax, where adding a new token always reduces the attention probability of existing tokens strictly, $\alpha$-entmax allows a distinct behavior: the attention probability can remain unchanged. This occurs because the $\alpha$-entmax’s thresholding effect allows tokens with logits below a certain threshold to receive exactly zero attention, letting the model focus only on the relevant tokens.
@@ 3.2 Attention Dispersion and Concentration
    This is impossible for $\alpha=1$ (softmax), for which we always have $\mathrm{softmax}(\bm{z})_{i}<\mathrm{softmax}(\bm{z}^{*})_{i}$.
@@
    $\lim_{n\to\infty}\frac{H(\mathrm{softmax}(\bm{z}_{1:n}/\theta))}{\log n}=1.$ (6)
@@
The full proof can be consulted in §C. The key takeaway of this result is that the entropy of attention distributions reveals how concentrated or dispersed they are across tokens. While softmax distributions approach maximum entropy $\Theta(\log n)$ as the sequence length increases (indicating complete dispersion), $\alpha$-entmax distributions maintain bounded entropy $\mathcal{O}(\log s)$ where $s$ is the support size. This allows models with $\alpha$-entmax to maintain focused, low-entropy attention patterns even when processing extremely long sequences, as long as the support size is smaller than the full sequence length. This non-dispersion property means transformers with $\alpha$-entmax can scale to very long contexts without the attention becoming dispersed, maintaining their ability to focus on relevant information regardless of how much additional context is present. However, attention dispersion is not the only obstacle to effective long-sequence modeling.
@@ 3.3 Representational Preservation and Over-squashing Alleviation
Two other critical challenges in long-context transformers are representational collapse and over-squashing, both exacerbated by the diffuseness of softmax attention (Barbero et al., 2024). More concretely, collapse means $\|\bm{v}_{i}^{(L)}-\bm{v}_{j}^{(L)}\|_{1}\!\to\!0$ as $n\!\to\!\infty$, and over-squashing means gradients from distant inputs vanish across $\mathcal{O}(n^{L})$ paths. Here, we show that $\alpha$-entmax mitigates both via sparse, bounded-support attention:
@@
[Representational Preservation and Reduced Gradient Paths] Let $\alpha>1$. Consider a depth-$L$ transformer with residual connections and attention weights given by $\alpha$-entmax. Suppose each attention distribution has support size at most $s$ with $s\ll n$. Then:
  1. 1.
    (Preserved representations) There exist input families $\bm{v}^{(0)}\!\in\!\mathbb{R}^{n\times d}$ and $\bm{v}^{*(0)}\!\in\!\mathbb{R}^{(n+1)\times d}$ and a constant $c>0$ such that $\|\bm{v}_{n}^{(L)}-\bm{v}^{*(L)}_{n+1}\|_{1}\geq c$ for all $n$; i.e., $\alpha$-entmax can maintain distinct token representations as $n\to\infty$.
  2. 2.
    (Alleviated over-squashing) The number of effective gradient paths scales as $\mathcal{O}(s^{L})$ (rather than $\mathcal{O}(n^{L})$ under softmax), alleviating over-squashing via stronger gradient signals.
Full statements and proofs are in App. §D.2–§D.3.
Empirical evidence.
We corroborate Proposition 3.3 with controlled experiments. For representational collapse, following the Barbero et al. (2024)-style one-token extension probe, we implement a counterexample with $\alpha\in\{1.0,1.5,1.75,2.0\}$, and report the $L_{1}$ gap between $\bm{v}_{n}^{(L)}$ and $\bm{v}_{n+1}^{*(L)}$. With softmax attention, the gap rapidly decays toward 0 with length, while $\alpha$-entmax preserves a non-vanishing margin up to 128K tokens, with larger $\alpha$ yielding stronger preservation (App. §D.2). For over-squashing, we analyzed gradient flow through an 8-layer network on a copying task, where the model must copy information across long distances. We find that $\alpha$-entmax attention sustains substantially larger norms across depths and lengths, consistent with $\mathcal{O}(s^{L})$ path growth, whereas softmax degrades sharply with sequence length (App. §D.3).
@@ 4.1 Controlling Sparsity in Long Contexts via ASEntmax
As sequence length grows, the spread of attention logits increases—for IID Gaussian logits, the expected range satisfies $\mathbb{E}[\Delta]\sim 2\sigma\sqrt{2\log n}$ (Kamath, 2015). With a fixed temperature, this makes attention overly peaky at long $n$. We address this with Adaptive-Scalable $\alpha$-entmax (ASEntmax), which rescales logits per head as a function of context length and content:
@@
where $\bm{w}_{\beta},\bm{w}_{\gamma}\in\mathbb{R}^{d}$ are learnable, head-specific projection vectors. This characterization allows the model to learn a slowly rising ($\gamma>0$) or dampening ($\gamma<0$) temperature schedule without interfering in the positional encodings (which would happen with negative values of $\beta$). Specifically, for IID Gaussian logits $\mathcal{N}(0,\sigma)$, when $\delta=0$ and $\gamma=-0.5$ the scaling counteracts the growth of logit ranges ($\Delta_{n}$) as context increases:
@@
Figure 3: Learned positions per head. Besides a simple linear fit baseline ($\beta n$), we also show the fit given by $\delta+\beta\log n$ and $\delta+\beta(\log n)^{\gamma}$, which are used by SSMax and ASEntmax, respectively.
@@
To empirically validate the importance of the parameter $\gamma$ in our proposed scaling formulation, we conducted experiments on a language modeling task using a 120M-parameter transformer trained on 5B tokens from the FineWeb dataset (Penedo et al., 2024). Following the methodology from SSMax (Nakanishi, 2025), we implemented learnable scaling parameters for the attention logits, but with a key difference: while Nakanishi (2025) uses a global scaling parameter, we learn separate scaling parameters for each attention head, motivated by Correia et al. (2019)’s finding that attention heads develop distinct sparsity patterns. Figure 3 presents the learned scaling behaviors for representative attention heads, along with fitted curves from different scaling models. First, we note that different heads learn significantly distinct patterns, highlighting the need of head-specific scaling. Second, the results demonstrate that a simple log-scaling model $\delta+\beta\log n$ provides poor fits for many heads. In contrast, the inclusion of a $\gamma$ power provides consistently better fits across different attention heads. The complete distribution of fitted $\beta$ and $\gamma$ values across all heads is provided in §F, and additional training details can be found in §F.1.
@@ 5.1 Synthetic Tasks
We use small decoder-only transformers and keep the number of layers as low as possible so that the results reflect the attention method’s capabilities rather than model scale. The only exception is Reverse, where we raise $L$ until a plain softmax baseline reaches at least $1.5\times$ the in-distribution performance. For positional information we default to NoPE+ALiBi, which we dub as NAPE, where half the heads have no positional encoding (NoPE) and the other half use ALiBi with linear slopes. [^1] Hard-ALiBi (Jelassi et al., 2024) allows zero slopes on some heads, which is equivalent to NoPE; our NAPE default mirrors this practical configuration. We treat NAPE as a practical default, not a contribution, and report RoPE and standalone ALiBi in App. §H for completeness. [^2] Across all attention methods tested—including softmax—NAPE consistently performs best. We study its interaction with $\alpha$-entmax and positional encodings in §E, with further empirical insights in §E.4 and §H.9. SSMax and ASEntmax models use $\delta=1$. Further hyperparameters and ablations appear in App. §G.